1. (10 points) Find an equation of the line of intersection of the planes 2 + 2y +32 = 2 2 + y + z = 1
Answer #1

The equation of two planes are x+y+z=1 x+2y+32= 2 — eleminating a from ① and 4 substracting 0 from ① we get x+2y +3Z-x-y-z= 2